OpenStudy (anonymous):

50. One solution of the equation 3x2 - 16x + 5=0 is 1/3. What is another solution ?

OpenStudy (anonymous):

x=1/3 x=5

OpenStudy (radar):

1/3 3 -16 5 1 -5 ---------------------- 3 -15 0 3x-15=0 x=5 Using synthetic division (dividing by x-1/3)